OpenStudy (anonymous):

HELLLLPPP DOUBLE MACBURGER PLEASE :) solve the given equation.check your solution 3|2a+7|=3a+12

OpenStudy (anonymous):

I got the answer as a = -3 or a = -11/3 . O_O

OpenStudy (anonymous):

yayyy!! that's one of my answer choices. Thank you :)

OpenStudy (anonymous):

Your welcome. =]

OpenStudy (anonymous):

*You're =P

OpenStudy (jiteshmeghwal9):

|a|=a &-a 3(2a+7)=3a+12 & 3*-1(2a+7)=3a+12

OpenStudy (jiteshmeghwal9):

solve this two equations :)